Train A has a speed 5 miles per hour greater than that of train B. Train A travels 195 miles in the same time train B travels 180 miles.
Form the equation that describes the problem.

Answer by mananth(16949)

180/x = 195/(x+5)
180(x+5)=195x
180x+900=195x
15x=900
x=60
Train B speed = 60 mph
Train A speed = 65 mph
CHECK
180/60 = 195/65
3=3